OpenStudy (anonymous):

PR has endpoints P(12,6) and R(-8,18). Find the length of PR to the nearest tenth. Will give medal to correct answer. :)

OpenStudy (jdoe0001):

\(\large \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (anonymous):

I have that part. I got to step 3 and got stumped.

OpenStudy (jdoe0001):

\(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \quad \\ P(12,6) \quad and \quad R(-8,18)\\ \quad \\ d = \sqrt{((-8)-(12))^2 + ((18)-(6))^2}\\\implies d = \sqrt{20^2+12^2} \implies d = \sqrt{400+144}\)

OpenStudy (jdoe0001):

well... to be exact....\(\bf d = \sqrt{((-8)-(12))^2 + ((18)-(6))^2}\\\implies d = \sqrt{(-20)^2+12^2} \implies d = \sqrt{400+144}\)

OpenStudy (anonymous):

That makes sense, but when I try to finish the equation, I don't get the right answer. :/ Like, it doesn't look like it is supposed to.

OpenStudy (jdoe0001):

hmmm what are your choices?

OpenStudy (anonymous):

They didn't give me any for this question. :/

OpenStudy (jdoe0001):

hmm... well... ahemm.... that's about the PR length

OpenStudy (anonymous):

Would it be -8?

OpenStudy (anonymous):

Nom wait...Would the answer be 23.33?

OpenStudy (anonymous):

No* not nom. lol

OpenStudy (jdoe0001):

no, is a distance amount between 2 points, thus the value will be positive, but not -8

OpenStudy (jdoe0001):

\(\bf \sqrt{400+144}\implies \sqrt{544}\approx 23.32380758\)

OpenStudy (anonymous):

So 23.33 or 23.32?

OpenStudy (jdoe0001):

well, if you want to round it up to 2 decimals, 23.32 notice after the "2" is a 3, not a 5, thus the "2" doesn't round up to 3 :)

OpenStudy (anonymous):

You kind of lost me with that last comment. lol

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2 \) a = 20; b = 12 \(20^2 + 12^2 = c^2 \) \(c^2 = 400 + 144\) \(c = \sqrt{544} \)